# Numerical simulation of type II primordial black hole formation

### v3

@@ 3.2 Identification of BH apparent horizon
First, let us introduce two independent future-directed radial null vector fields, $l$ and $l^{\prime}$. We also define the null expansion for $l$ and $l^{\prime}$ as $\theta$ and $\theta^{\prime}$. A null expansion is expressed as $\theta\coloneqq 2\mathcal{L}_{l}R/R$, where $\mathcal{L}_{l}$ denotes the Lie derivative along $l$. Trapping horizons are defined by the surfaces that satisfy $\theta\theta^{\prime}=0$ or equivalently the compactness $12M/R=$ 1 with $M$ being the Misner–Sharp mass defined by $M(t,r)\coloneqq R(1-\nabla_{\mu}R\,\nabla^{\mu}R)/2$ [29, 41, 42, 8].
In Fig. 5, the snapshots of the compactness $\quantity(2M/R)(z)$ are plotted. For $\mu=0.5$, the intersection between $\quantity(2M/R)(r)$ and the horizontal line of $12M/R=$ 1 is always only one. On the other hand, for the cases $\mu=1.2$ and $1.8$, the line intersects the $12M/R=$ 1 more than three times at a sufficiently late time. The worldlines of the intersections are shown in Fig. 6. To identify the future-trapped ($\theta<0$ and $\theta^{\prime}<0$) and past-trapped ($\theta>0$ and $\theta^{\prime}>0$) regions we also show directions of two independent null vector fields $l$ and $l^{\prime}$ together with the contour curves of the areal radius in Figs. 7 and 8. Comparing the directions of the null vector fields and contour curves of the areal radius, we can realize that the red and blue shaded regions in Fig. 6 are future- and past-trapped regions, respectively. More specifically, in the future(past)-trapped region, both arrows direct to directions along which the areal radius decreases (increases), and the curve described by red points indicates the boundary satisfying $\theta\theta^{\prime}=0$. Then, we refer to a marginally future trapped surface that is the outer boundary of the future-trapped region on the $t=$const hypersurface as a BH apparent horizon. Similarly, we refer to a marginally past-trapped surface that is the inner boundary of the past-trapped region on the $t=$ const hypersurface as a cosmological apparent horizon.
It is not easy to understand the horizon configuration. Therefore, we carefully investigate it and present a more concrete classification of trapping horizons in the following sections. Readers who would like to get a quick intuition about the horizon configuration may refer to the schematic conformal diagram in Figures 16 and 17.
@@
Figure 7: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted as orange and blue arrows for $\mu=1.2$ case (type I) in the middle panel of Fig. 6. $R$-contour with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top is plotted as black curves.
Figure 8: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted as orange and blue arrows for $\mu=1.8$ case (type II) in the right panel of Fig. 6. The $R$ contour is plotted as black curves with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top. In addition to the original plot of $R$-levels with 1.0 increments, it is also plotted with 0.1 increments from $R=8.1$ to $8.9$.
3.3 Spacetime structures of type I/II PBH
Let us introduce the classification of trapping horizons following Refs. [42, 43]. For this purpose, to avoid confusion, we introduce another notation for the two radial null vector fields and associated expansions as $l_{\pm}$ and $\theta_{\pm}$ for a specific trapping horizon. Then we always set $\theta_{+}=0$ when considering a trapping horizon satisfying $\theta_{+}\theta_{-}=0$. The classification of trapping horizons can be summarized as Table 1.
@@ 3.3 Spacetime structures of type I/II PBH
Table 1: Classification of trapping horizons satisfying $\theta_{+}\theta_{-}=0$ [42]. A black hole, white hole, and cosmological trapping horizons with $w=1/3$ $(w=0)$ are identified with a future-outer, past-outer, and past-degenerate (past-inner) trapping horizons, respectively.
(a) Type I PBH
(b) Type II PBH
Figure 9: The conformal diagrams of type I and II PBHs in radiation domination inferred from the present numerical results. The left figure defines a type I PBH as a causal structure of the typical case of PBH. The right one includes bifurcating trapping horizons and defines a type II PBH. The red and blue curves depict the future and past trapping horizons, respectively. The bifurcating trapping horizons ($\theta_{+}=\theta_{-}=0$) are depicted as circles in the type II case.
Unlike the dust case discussed in Appendix A, we can only obtain numerical solutions such as those shown in Figs. 7 and 8 for the radiation case. From these figures and carefully comparing the similarities and differences from the exact conformal diagrams of the dust solution discussed in Appendix A, we can deduce possible conformal diagrams and proper identifications of the trapping horizons for the radiation case as shown in the left and right panels in Fig. 9 as possible conformal diagrams of type I and type II PBHs, respectively.
The crucial difference between the right and left panels in Fig. 9 is the existence of the bifurcating trapping horizons $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(0;*)$, which is described by the intersection points of the two trapping horizon trajectories in Fig. 8. In this paper, we propose to use this difference to classify PBH formation processes into types I and II; that is, we identify the PBH as of type I if there is no bifurcating trapping horizon and of type II if there appears at least one bifurcating trapping horizon during the time evolution associated with the intersection point. This difference has already been pointed out by KHW in Ref. [1] for the case of the dust fluid. Unlike the current proposal, the authors in Ref. [1] did not clearly distinguish the definitions of type I/II PBHs and type I/II fluctuations. However, in our simulation with $w=1/3$, we find that the horizon configuration is of type I for the $\mu=1.6$ case, but the initial fluctuation is of type II. Therefore, there is a gap between the definitions of type I/II PBHs and type I/II fluctuations. This fact is our motivation for the new definition of the PBH.
@@
It is worth noting that the schematic diagram given by Fig. 9 implies that all light rays emitted from the center reach not the future null infinity but the future singularity in the type II PBH. Although we cannot make any general statement for the moment, the null trajectories indicated by the orange arrows in Fig. 8 seem consistent with this observation.
Precisely, one could also consider the critical case between type I and type II PBHs when the two bifurcating trapping horizons degenerate, then we additionally classify this case as a type IId PBH, which has just one bifurcating trapping horizon.
@@ 4.1 PBH mass with the Gaussian initial fluctuation
where $\rho_{b}=\rho_{b,0}(t_{0}/t)^{2}$ is the background energy density, and $F$ is a constant that corresponds to the efficiency of accretion, which is commonly numerically found to be of order $\mathcal{O}(1)$. It is important to note that Eq. (4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion as Carr and Hawking (1974) [4] already pointed out. However, it will correctly describe the accretion into the PBHs for sufficiently large times once we have a quasi-stationary flow into the PBH. The integrated solution is,
@@
Figure 11: PBH mass in terms of the horizon mass at horizon entry $M_{H}$ to the amplitude of fluctuation $\mu$. In this case, in terms of fluctuation types, $\mu=1.2$ is classified as type I, while $\mu=1.4\>\text{to}\>2.2$ as type II, as seen in Fig. 2. In terms of formation types, on the other hand, $\mu=1.2$ to $1.6$ are classified as type I, while $\mu=1.8$ to $2.2$ as type II, as seen in Figs. 6 and 9.
The initial and final PBH masses get larger with increasing amplitude $\mu$. We will check the profile dependence in the next subsection.
@@ 4.2 PBH mass with the initial fluctuation motivated by the three-zone model
In order to construct a specific initial profile that can be contrasted with the former Gaussian profile, let us try to utilize the three-zone profile as a reference, which is an extension of the model used in Ref. [1].
To realize a similar situation to the schematic conformal diagram, let us consider the closed FLRW region for the initial overdense region. Then, as is shown in the Appendix B, under the boundary condition $\zeta=0$ at the surface of the overdense region, we obtain the following form of $\zeta(r)$:
@@
The behavior of the horizon configuration is similar to the case of the Gaussian profile, and we do not explicitly show it. In contrast to the Gaussian profile, the initial PBH mass formed by this profile decreases as the amplitude of fluctuation $\tilde{\mu}$ increases for $\tilde{\mu}>0.65$ as shown in Fig. 14, even though they are classified as type I fluctuations and formations for $\tilde{\mu}\leq 0.88$. This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases with the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of non-linear perturbation. We also find the gap between the parameter region of type I/II fluctuations and type I/II PBHs, as indicated in Fig. 14.
@@ 5 Conclusion
We have numerically investigated the PBH formation from a type II initial fluctuation $\zeta(r)$, which is defined by the non-monotonicity of the areal radius $R(r)\propto re^{\zeta}$ with respect to the coordinate radius $r$ on the spatial geometry, i.e., the existence of $\partial_{r}R(r)=0$ (called neck or throat structure), in the radiation-dominated situation. As in the dust case reported in KHW, the PBH formation process can be classified into types I and II according to the configuration of the trapped regions $R(t,r)\leq 2M(t,r)$ (or trapping horizons $R(t,r)=2M(t,r)$). Specifically, if there is a bifurcating trapping horizon, the PBH is classified into type II, and type I otherwise. In the dust case, KHW reported that a type I fluctuation results in a type I PBH, and a type II fluctuation does in a type II PBH. However, in the case of radiation fluids, this is not true. We found a gap between the classifications regarding the shape of the initial spatial geometry (fluctuation type) and the structure of trapping horizons (formation type). That is, we found that a type II fluctuation may result in a type I PBH. Therefore, we propose classifying PBHs based on the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
In the leading order of the long-wavelength limit, the overdense region can be roughly approximated by the closed FLRW universe, and the amplitude of the density perturbation can be translated into the value of the radial coordinate at the edge of the closed FLRW region. Then, the type II fluctuation indicates that the edge is located in the other hemisphere of $S^{3}$ initial hypersurface beyond the great circle. Therefore, in the case of type II PBHs, one may expect that the PBH mass would decrease with increasing the amplitude of initial fluctuation since the initial time slice of $S^{3}$ in the overdensity region closes in the limit of large-density perturbation, which indicates the area of the edge is vanishing. However, this is not the case for the Gaussian-type initial fluctuation profile, while the other profile exhibits decreasing mass with increasing initial amplitude. Therefore, we conclude that the behavior of the initial PBH mass depends on not only the type of fluctuation but also the fluctuation profile. Thus, systematic investigation of the profile dependence would be needed, e.g., by using the shape parameter defined as the curvature peak of the compaction function $\mathcal{C}_{\text{SS}}$ [19, 45]. In addition to the profile dependence, the dependence on the equation of state $p=w\rho$ is also an interesting issue, which fills the gap between KHW and this work [20].
@@ Acknowledgments
The authors are grateful to K.-I. Nakao, B.J. Carr, T. Igata, and J.M.M. Senovilla for useful comments and fruitful discussion. A.E. acknowledges support from the JSPS Postdoctoral Fellowships for Research in Japan (Graduate School of Sciences, Nagoya University). T.H. is supported in part by JSPS KAKENHI Grant Nos. 19K03876 and 20H05853. D.S. would like to take this opportunity to thank the “Nagoya University Interdisciplinary Frontier Fellowship” supported by Nagoya University and JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2120. C.Y. is supported in part by JSPS KAKENHI Grant Nos. 20H05850 and 20H05853. C.Y. and T.H. acknowledge the hospitality at APCTP during the focus research program ”Black Hole and Gravitational Waves: from modified theories of gravity to data analysis”, where part of this work was discussed.
@@ Appendix A Conformal diagrams of the three-zone model for the dust case
In the three-zone model [7], the spacetime of PBH is described by the patchwork of the three different regions: the innermost overdense region is given by a spatially closed FLRW universe, the intermediate underdense region, and the outermost flat FLRW universe as the background. We also note that the intermediate underdense region must be introduced to compensate for the mass excess of the innermost region and connect the spacetime to the background spacetime without a mass gap. Without the intermediate underdense region, this model cannot be the exact solution, even for the dust case.
@@
For the underdense region with a perfect fluid with $w=1/3$, it would not be tractable to obtain any analytical and exact solutions of such dynamical inhomogeneous spacetimes with non-linear sound waves. Therefore, we stick to the three-zone model with dust. Furthermore, for simplicity, we focus on the model where the underdense region is described by the Schwarzschild solution, where the matching is possible without any singular hypersurface for the dust case. Despite this great simplification, the model provides a very powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons. Note that one can also approach the same problem by using the Lemaître-Tolman-Bondi exact dust solution (c.f. [1]).
@@
The conformal diagrams for the three spacetimes of which the three-zone model consists are shown in Fig. 15, where future $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;*)$, past (+;)(+;*)( + ; ∗ ), and bifurcating $(0;*)$ trapping horizons are denoted by red lines, blue lines, and the intersection point between them in the Schwarzschild region, respectively. For the dust case $w=0$, the BH apparent horizon and cosmological apparent horizon become future-outer (;)(-;-)( - ; - ) and past-inner (+;+)(+;+)( + ; + ), denoted by red -solid and blue-dashed lines, respectively.
@@
First, let us construct the spacetime diagram of the case of the type I fluctuation. The extracted regions for the patchwork in the type I case are described by the yellow shaded regions in Fig. 15, and the resultant spacetime diagram is shown in the left panel of Fig. 17. These conformal diagrams define the right and left-directed future null vectors as $l_{\rm out}$ and $l_{\rm in}$, respectively. For the closed FLRW region, on the boundary of the extracted region, we can find $\theta_{\rm out}>0$ and $\theta_{\rm in}<0$ at the maximum expansion. Therefore, in the outside Schwarzschild region, we also require the same sign of $\theta_{\rm out}$ and $\theta_{\rm in}$. Similarly, one can understand the matching of the Schwarzsdhild region and the background spatially flat FLRW spacetime.
@@
The extracted regions are yellow-shaded for the type II fluctuation in Fig. 16. Since the type II fluctuation should have a maximal value of the areal radius, the boundary of the closed FLRW region must be outside the central point at which the areal radius takes the maximum value corresponding to the area of the big sphere at the maximum expansion. Then we find $\theta_{\rm out}<0$ and $\theta_{\rm in}>0$ on the boundary, which can be matched with the Schwarzschild region only in the left part of the Schwarzschild conformal diagram. Thus, the resultant conformal diagram for type II is given by the right panel of Fig. 17.
@@
Figure 15: Conformal diagrams to use for a patchwork for the spacetime of the type I PBH schematically in Fig. 17 for dust case $w=0$. Dashed gray curves represent $R$-levels. Red(blue)-shaded regions represent future(past) trapped regions. Future-outer $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;-)$, past-outer (+;)(+;-)( + ; - ), future-inner (;+)(-;+)( - ; + ), and past-inner (+;+)(+;+)( + ; + ) trapping horizons are denoted by red-solid, blue -solid, red -dashed, and blue -dashed lines, respectively.
@@
Figure 16: Same as Fig. 15 but for the type II PBH. We use the yellow-shaded regions for a patchwork.
@@
Figure 17: Conformal diagrams of PBH for dust case $w=0$ as a patchwork of the three diagrams and trapping horizons. The left figure defines a type I PBH as a typical case PBH causal structure. The right one is constructed by changing the cutting diagram range and defining a type II PBH.

### v4

@@ 3.2 Identification of BH apparent horizon
First, we introduce two independent future-directed radial null vector fields, $l$ and $l^{\prime}$. We also define the null expansion for $l$ and $l^{\prime}$ as $\theta$ and $\theta^{\prime}$. A null expansion is expressed as $\theta\coloneqq 2\mathcal{L}_{l}R/R$, where $\mathcal{L}_{l}$ denotes the Lie derivative along $l$. Trapping horizons are defined by the surfaces that satisfy $\theta\theta^{\prime}=0$ or equivalently the compactness $12M/R=$ 1 with $M$ being the Misner–Sharp mass defined by $M(t,r)\coloneqq R(1-\nabla_{\mu}R\,\nabla^{\mu}R)/2$ [29, 41, 42, 8].
In Fig. 5, the snapshots of the compactness $\quantity(2M/R)(z)$ are plotted. For $\mu=0.5$, the intersection between $\quantity(2M/R)(r)$ and the horizontal line of $12M/R=$ 1 is always only one. On the other hand, for the cases $\mu=1.2$ and $1.8$, the line intersects $12M/R=$ 1 more than three times at a sufficiently late time. The worldlines of the intersections are shown in Fig. 6. To identify future-trapped ($\theta<0$ and $\theta^{\prime}<0$) and past-trapped ($\theta>0$ and $\theta^{\prime}>0$) regions, we also show directions of two independent null vector fields $l$ and $l^{\prime}$ together with the contour curves of the areal radius in Figs. 7 and 8. Comparing the directions of the null vector fields and the contour curves of the areal radius, we can realize that the red and blue shaded regions in Fig. 6 are future- and past-trapped regions, respectively. More specifically, in the future(past)-trapped region, both arrows direct to directions along which the areal radius decreases (increases), and the curve described by red points indicates the boundary satisfying $\theta\theta^{\prime}=0$. Then, we refer to a marginally future-trapped surface that is the outer boundary of the future-trapped region on the $t=$const hypersurface as a BH apparent horizon. Similarly, we refer to a marginally past-trapped surface that is the inner boundary of the past-trapped region on the $t=$ const hypersurface as a cosmological apparent horizon.
It is not easy to understand the horizon configuration. Therefore, we carefully investigate it and present a more concrete classification of trapping horizons in the following sections. Readers who would like to get a quick intuition about the horizon configuration can refer to the schematic conformal diagram in Figures 16 and 17.
@@
Figure 7: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted as orange and blue arrows for $\mu=1.2$ case (type I) in the middle panel of Fig. 6. The $R$ contour is plotted as black curves with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top.
Figure 8: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted as orange and blue arrows for $\mu=1.8$ case (type II) in the right panel of Fig. 6. The $R$ contour is plotted as black curves with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top. In addition to the original plot of $R$ levels with increments of 1.0, it is also plotted with increments of 0.1 from $R=8.1$ to $8.9$.
3.3 Spacetime structures of type A/B PBH
Let us introduce the classification of trapping horizons following Refs. [42, 43]. For this purpose, to avoid confusion, we introduce another notation for the two radial null vector fields and associated expansions as $l_{\pm}$ and $\theta_{\pm}$ for a specific trapping horizon. Then we always set $\theta_{+}=0$ when considering a trapping horizon satisfying $\theta_{+}\theta_{-}=0$. The classification of the trapping horizons can be summarized in Table 1.
@@ 3.3 Spacetime structures of type A/B PBH
Table 1: Classification of trapping horizons satisfying $\theta_{+}\theta_{-}=0$ [42]. A black hole, a white hole, and cosmological trapping horizons with $w=1/3$ $(w=0)$ are identified with future-outer, past-outer, and past-degenerate (past-inner) trapping horizons, respectively.
(a) Type IA PBH
(b) Type IIB PBH
Figure 9: The present numerical results are inferred from the conformal diagrams of type A and B PBHs in radiation domination. The left figure defines a type A PBH as a causal structure of the typical case of PBH. The right one includes bifurcating trapping horizons and defines a type B PBH. The red and blue curves depict the future and past trapping horizons, respectively. The bifurcating trapping horizons ($\theta_{+}=\theta_{-}=0$) are depicted as circles in the type B case.
Unlike the dust case discussed in Appendix A, we can only obtain numerical solutions such as those shown in Figs. 7 and 8 for the radiation case. From these figures and carefully comparing the similarities and differences from the exact conformal diagrams of the dust solution discussed in the Appendix A, we can deduce possible conformal diagrams and proper identifications of the trapping horizons for the radiation case as shown in the left and right panels in Fig. 9 as possible conformal diagrams of type A and type B PBHs, respectively.
The crucial difference between the right and left panels in Fig. 9 is the existence of the bifurcating trapping horizons $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(0;*)$, which is described by the intersection points of the two trapping horizon trajectories in Fig. 8. In this paper, we propose to use this difference to classify the PBH formation processes into types A and B; that is, we identify the PBH as type A if there is no bifurcating trapping horizon and as type B if at least one bifurcating trapping horizon appears during the time evolution associated with the intersection point. KHW has already pointed out this difference in Ref. [1] for the case of the dust fluid. Unlike the current proposal, the authors in Ref. [1] identified the definitions of type I/II and type A/B. However, in our simulation with $w=1/3$, we find the case of type IIA, the horizon configuration is of type A from type II fluctuation for the $\mu=1.6$ case. This fact is our motivation for the new definition of the PBH.
@@
It is worth noting that the schematic diagram given by Fig. 9 implies that all light rays emitted from the center reach not the future null infinity but the future singularity of the type B PBH. Although we cannot make any general statement for the moment, the null trajectories indicated by the orange arrows in Fig. 8 seem consistent with this observation.
@@ 4.1 PBH mass with the Gaussian initial fluctuation
where $\rho_{b}=\rho_{b,0}(t_{0}/t)^{2}$ is the background energy density, and $F$ is a constant that corresponds to the efficiency of accretion, which is commonly numerically found to be of order $\mathcal{O}(1)$. It is important to note that Eq. (4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion as Carr and Hawking (1974) [4] already pointed out. However, once we have a quasi-stationary flow into the PBH, it will correctly describe the accretion into the PBHs for sufficiently long times. The integrated solution is,
@@
Figure 11: PBH mass in terms of the horizon mass at horizon entry $M_{H}$ to the amplitude of fluctuation $\mu$. In this case, in terms of fluctuation types, $\mu=1.2$ is classified as type I, while $\mu=1.4\>\text{to}\>2.2$ as type II, as seen in Fig. 2. In terms of formation types, on the other hand, $\mu=1.2$ to $1.6$ are classified as type A, while $\mu=1.8$ to $2.2$ as type B, as seen in Figs. 6 and 9.
The initial and final PBH masses get larger with increasing amplitude $\mu$. We will check the dependence of the profile in the next subsection.
@@ 4.2 PBH mass with the initial fluctuation motivated by the three-zone model
To construct a specific initial profile that can be contrasted with the former Gaussian profile, let us try to utilize the three-zone profile as a reference, an extension of the model used in Ref. [1].
Let us consider the closed FLRW region for the initial overdense region to realize a situation similar to the schematic conformal diagram. Then, as is shown in the Appendix B, under the boundary condition $\zeta=0$ at the surface of the overdense region, we obtain the following form of $\zeta(r)$:
@@
The behavior of the horizon configuration is similar to the case of the Gaussian profile, and we do not explicitly show it. In contrast to the Gaussian profile, the initial PBH mass formed by this profile decreases as the amplitude of fluctuation $\tilde{\mu}$ increases for $\tilde{\mu}>0.65$ as shown in Fig. 14, even though they are classified as type I fluctuations and formations for $\tilde{\mu}\leq 0.88$. This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases with the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of non-linear perturbation. We also find the gap between the parameter region of type I/II and type A/B PBHs, as indicated in Fig. 14.
@@ 5 Conclusion
We have numerically investigated the type II PBH formation from a type II initial fluctuation $\zeta(r)$, which is defined by the non-monotonicity of the areal radius $R(r)\propto re^{\zeta}$ with respect to the coordinate radius $r$ on the spatial geometry, i.e., the existence of $\partial_{r}R(r)=0$ (called neck or throat structure), in the radiation-dominated situation. As in the dust case reported in KHW, the PBH formation process can be classified into types I and II according to the configuration of the trapped regions $R(t,r)\leq 2M(t,r)$ (or trapping horizons $R(t,r)=2M(t,r)$). Specifically, if there is a bifurcating trapping horizon, the PBH is classified as type B, and otherwise, it is classified as type A. In the dust case, KHW reported that a type I fluctuation results in a type A PBH and a type II fluctuation in a type B PBH. However, in the case of radiation fluids, this is not true. We found a gap between the classifications regarding the shape of the initial spatial geometry (fluctuation type) and the structure of the trapping horizons (formation type). We found that a type II fluctuation may result in a type I PBH, so-called type II-A PBH. Therefore, we propose classifying type A/B PBHs based on the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
In the leading order of the long-wavelength limit, the overdense region can be roughly approximated by the closed FLRW universe, and the amplitude of the density perturbation can be translated into the value of the radial coordinate at the edge of the closed FLRW region. Then, the type II fluctuation indicates that the edge is located in the other hemisphere of $S^{3}$ initial hypersurface beyond the great circle. Therefore, in the case of type II PBHs, one may expect that the PBH mass would decrease with increasing the amplitude of initial fluctuation since the initial time slice of $S^{3}$ in the overdensity region closes in the limit of large-density perturbation, which indicates that the area of the edge is vanishing. However, this is not the case for the Gaussian-type initial fluctuation profile, while the other profile exhibits decreasing mass with increasing initial amplitude. Therefore, we conclude that the behavior of the initial PBH mass depends on not only the type of fluctuation but also the fluctuation profile. Thus, a systematic investigation of the profile dependence would be needed, e.g., by using the shape parameter defined as the curvature peak of the compaction function $\mathcal{C}_{\text{SS}}$ [19, 45]. In addition to profile dependence, dependence on the equation of state $p=w\rho$ is also an interesting issue, filling the gap between KHW and this work [20].
@@ Acknowledgments
The authors are grateful to K.-I. Nakao, B.J. Carr, T. Igata, and J.M.M. Senovilla for useful comments and fruitful discussion. A.E. acknowledges support from the JSPS Postdoctoral Fellowships for Research in Japan (Graduate School of Sciences, Nagoya University). The authors are supported in part by the JSPS KAKENHI Grant Nos. 19K03876(TH), 20H05850(CY), 20H05853(TH, CY), and 24KJ1223(DS). KU would like to take this opportunity to thank the “Nagoya University Interdisciplinary Frontier Fellowship” supported by Nagoya University and JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2125. CY and TH acknowledge the hospitality at APCTP during the focus research program ”Black Hole and Gravitational Waves: from modified theories of gravity to data analysis”, where part of this work was discussed.
@@ Appendix A Conformal diagrams of the three-zone model for the dust case
In the three-zone model [7], the spacetime of PBH is described by the patchwork of the three different regions: the innermost overdense region is given by a spatially closed FLRW universe, the intermediate underdense region, and the outermost flat FLRW universe as the background. We also note that the intermediate underdense region must be introduced to compensate for the innermost region’s mass excess and connect the spacetime to the background spacetime without a mass gap. Without the intermediate underdense region, this model cannot be the exact solution, even for the dust case.
@@
For the underdense region with a perfect fluid with $w=1/3$, obtaining any analytical and exact solutions of such dynamical inhomogeneous spacetimes with non-linear sound waves would not be tractable. Therefore, we stick to the three-zone model with dust. Furthermore, for simplicity, we focus on the model where the underdense region is described by the Schwarzschild solution, where the matching is possible without any singular hypersurface for the dust case. Despite this significant simplification, the model provides a powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons. Note that one can also approach the same problem by using the Lemaître-Tolman-Bondi exact dust solution (c.f. [1]).
@@
The conformal diagrams for the three spacetimes of which the three-zone model consists are shown in Fig. 15, where future $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;*)$, past (+;)(+;*)( + ; ∗ ), and bifurcating $(0;*)$ trapping horizons are denoted by red lines, blue lines, and the intersection point between them in the Schwarzschild region, respectively. For the dust case $w=0$, the apparent horizon and cosmological apparent horizon become future-outer (;)(-;-)( - ; - ) and past-inner (+;+)(+;+)( + ; + ), denoted by red-solid and blue-dashed lines, respectively.
@@
First, let us construct the spacetime diagram of the case of the type I fluctuation. The extracted regions for the patchwork in the type I (type A) case are described by the yellow shaded regions in Fig. 15, and the resultant spacetime diagram is shown in the left panel of Fig. 17. These conformal diagrams define the right and left-directed future null vectors as $l_{\rm out}$ and $l_{\rm in}$, respectively. For the closed FLRW region, on the boundary of the extracted region, we can find $\theta_{\rm out}>0$ and $\theta_{\rm in}<0$ at the maximum expansion. Therefore, in the outside Schwarzschild region, we also require the same sign of $\theta_{\rm out}$ and $\theta_{\rm in}$. Similarly, one can understand the matching of the Schwarzsdhild region and the background spatially flat FLRW spacetime.
@@
The extracted regions are yellow-shaded for the type II fluctuation in Fig. 16. Since the type II fluctuation should have a maximal value of the areal radius, the boundary of the closed FLRW region must be outside the central point at which the areal radius takes the maximum value corresponding to the area of the big sphere at the maximum expansion. Then we find $\theta_{\rm out}<0$ and $\theta_{\rm in}>0$ on the boundary, which can be matched with the Schwarzschild region only in the left part of the Schwarzschild conformal diagram. Thus, the resultant conformal diagram for type II (type B) is given by the right panel of Fig. 17.
@@
Figure 15: Conformal diagrams to use for a patchwork for the spacetime of the type I(type A) PBH schematically in Fig. 17 for dust case $w=0$. Dashed gray curves represent $R$-levels. Red(blue)-shaded regions represent future(past) trapped regions. Future-outer $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;-)$, past-outer (+;)(+;-)( + ; - ), future-inner (;+)(-;+)( - ; + ), and past-inner (+;+)(+;+)( + ; + ) trapping horizons are denoted by red-solid, blue-solid, red-dashed, and blue-dashed lines, respectively.
@@
Figure 16: Same as Fig. 15 but for the type II(type B) PBH. We use the yellow-shaded regions for a patchwork.
@@
Figure 17: Conformal diagrams of PBH for dust case $w=0$ as a patchwork of the three diagrams and trapping horizons. The left figure defines a type I(type A) PBH as a typical case PBH causal structure. The right one is constructed by changing the cutting diagram range and defining a type II(type B) PBH.